# Noise-based reward-modulated learning

### v3

@@ 1 Introduction
Backpropagation (BP) remains the workhorse algorithm for training artificial neural networks, computing precise global gradients for weight updates [46]. However, its operational requirements are fundamentally incompatible with the constraints of neuromorphic hardware [10]. BP requires precise global error propagation, which translates into high-bandwidth global data movement and large memory requirements to store activations, the main source of energy consumption in deep learning systems [36]. This contrasts sharply with the hardware needs for local synaptic updates that rely only on information available at the synaptic weight. In addition, BP’s iterative forward and backward passes are incompatible with the asynchronous event-driven regime in which most neuromorphic devices operate [41], and its deterministic nature cannot tolerate inherit noise in neuromorphic circuits.
In response, local plasticity approaches and biologically-inspired learning rules have been proposed to bypass the constraints of BP [64, 63, 32, 47, 24, 67]. A particularly promising family of methods is noise-based learning, which leverages stochastic perturbations (either in weights [12, 69, 5, 9] or activations [18, 23, 65, 62, 15]) to approximate the true direction of the gradient. By injecting small random perturbations, these methods avoid the need for perfectly matched feedback pathways and can leverage inherit noise [51, 35].
Given the prevalence of noise on neuromorphic and biological substrates [11, 53, 37], using noise as a mechanism to learn synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [33, 38] as a promising candidate. RMHL, which is a form of three-factor Hebbian learning [20, 30], offers a potentially powerful mechanism for credit assignment without explicit backpropagation of errors. Despite theoretical advances, few noise-based methods have been adapted to real-world tasks. Some attempts which have been made include application to control problems [4], for spatiotemporal pattern generation [27], and for training non-differentiable spiking neural networks training [14], and these particularly struggle with delayed reward regimes. This gap has limited the integration of noise-based approaches for effective credit assignment in temporally extended setups.
@@ 2.2.2 Noise-based learning rule
where $K$ denotes the number of samples and $\tilde{\theta}_{t}^{(i)}=\theta_{t}+\epsilon^{(i)}_{t}$ the noise-perturbed parameters (see Appendix A). In practice, we consider $K=1$, analogous to single-sample updates in stochastic gradient descent. Here, $g\bigl(\tilde{\theta}^{(i)}_{t}\bigr)=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})$ and $g\left(\theta_{t}\right)=\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$ are the log of the noise-perturbed and noise-free output, respectively. Thus, we define $\rho_{t}=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})-\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$, which captures the impact of the noise on the policy. Putting this together for the $K=1$ case, we obtain our noise-based learning rule
@@ 2.2.3 Learning with node level noise in neural network
To formulate the learning rule for specific layer parameters $W^{l}_{t}\in\{W^{1}_{t},\ldots,W^{L}_{t}\}$, we re-express the gradient in terms of the pre-activations instead of the parameters themselves. Here, the general parameter vector $\theta_{t}=\{W^{1}_{t},\ldots,W^{L}_{t}\}$ corresponds to the full set of all weight matrices for all layers at time $t$. To do so, we first rewrite the gradient term as
$\nabla g\left(W^{l}_{t}\right)=\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)$ (10)
and apply the chain rule to break down the gradient with respect to $W^{l}_{t}$ as
@@
where $i$ indexes over a set of repeated samples of the noise term, $\xi$.
Again, as above, by reducing this to the single sample ($n=1$) case, we can define a layer-specific weight update rule
@@ 2.2.4 Learning from sparse rewards
Our derivations so far assume that synaptic updates happen at every time step, implying a continuous stream of rewards and learning signals. In reality, however, rewards are often sparse, arriving only after a sequence of actions or upon reaching specific milestones. To handle this, we’ll modify our learning rule to account for rewards received at arbitrary times.
Let us denote these discrete reward times as $\tau_{m}\in\{\tau_{0},\tau_{1},...\tau_{M}\}$. At these moments, we compute the RPE as
@@ 2.4 Experimental validation
We compare NRL to two baselines: an exact-gradient version of NRL, which serves as an ”optimal performance” benchmark, and an RMHL approach.
@@
Table 1: Learning rules. The terms in the BPTT update are rearranged for easier comparison.
@@
NRL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{\,l}_{t}\,\rho_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$
BPTT $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\delta_{\tau_{m}}\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})$
RMHL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{\,l}_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$
In environments with high reward variability, like Cartpole and Acrobot, we stabilize synaptic updates by scaling the RPE by dividing by $r_{\tau_{m}}$. Hyperparameter values and training details are provided in Appendix B. All experiments are implemented in Python using PyTorch [43]. Our models and scripts are available for reproducibility at https://github.com/jesusgf96/noise-based-reward-modulated-learning.
@@ 3.1 Performance on control tasks with instantaneous and delayed rewards
To ensure comparability with RMHL methods, which typically utilize single-hidden-layer networks, we first conduct experiments with a one-hidden-layer neural network for BP, NRL, and RMHL. Training details are provided in Appendix B. In Section titled “Scalability to deeper architectures”, we extend the comparison to deeper networks with multiple hidden layers.
@@ 3.3 Learning using only noisy passes
To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem. The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Fig. 4A. Furthermore, Fig. 4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with a minimum of 2 noisy passes. We also extended this evaluation to 10 noisy passes to explore the impact of increasing the number of passes, showing slightly faster initial convergence.
@@ 4 Discussion
Our results demonstrate the learning effectiveness of NRL, solving both immediate and delayed feedback/reward problems in simulated environments, achieving performance comparable to baselines optimized employing backpropagation. In addition, it shows a significant improvement over RMHL [33, 38], a modern approach that also uses noise as the core of its synaptic plasticity. Our scalability results further demonstrate that NRL effectively assigns credit even in multilayer networks, a challenge for many biologically plausible or local algorithms, indicating its potential for scaling. This is a critical difference from RMHL, which struggles to assign credit in multilayer networks, as evidenced by its minimal performance improvements across trials. Nevertheless, as the depth of the network increases, NRL shows slower convergence due to its noise-driven updates.
NRL shares similarities with node perturbation methods [18, 23, 65, 62, 15], which inject noise into neurons to adjust weights based on state and performance fluctuations. However, unlike node perturbation, which requires constant feedback or reward for learning, NRL is well-suited for delayed reward setups. By leveraging eligibility traces [22, 25], it can effectively assign credit to past neural states even when actions and outcomes are separated in time.
@@
Additionally, the local-update structure of NRL provides direct compatibility with neuromorphic substrates [50], where only locally available information in the synapse is prioritized and global communication is severely penalized. This eliminates the need for the high-bandwidth global data movement and large memory buffers required by backpropagation, the primary source of energy consumption in deep learning models, addressing the fundamental energy and latency bottlenecks of conventional Von Neumann systems. Furthermore, NRL’s reliance on noise for gradient approximation offers a paradigm shift: it transforms the device stochasticity common in neuromorphic circuits from an engineering challenge into a computational resource. The inherent noise in NRL is also advantageous for reinforcement learning setups, as it naturally facilitates the exploration-exploitation balance [26, 31]. Injected noise perturbs neural states, promoting exploratory deviations from expected values and thus introducing variations in the agent’s policy, which benefits adaptation on neuromorphic edge devices.
@@
Despite its promise, our work has limitations that suggest future research directions. The current requirement for multiple forward passes (noisy and clean) per synaptic update, while theoretically simple, introduces a time-latency overhead. Future work should explore more efficient ways to estimate noise impacts, potentially eliminating the need for multiple passes. In addition, while we focused on a rate-based model, the deployment of NRL on physical hardware like Intel Loihi [8] requires its extension to spiking neural networks (SNNs). Furthermore, the simplicity of our RPE calculation limits its robustness in more complex tasks requiring long-term planning. Future development should incorporate more sophisticated temporal difference learning mechanisms [54, 56], such as a learned value function within an actor-critic architecture [28], to provide a richer temporally sensitive estimate of rewards. Finally, while backpropagation may offer faster convergence in unconstrained settings, recent studies [7, 13], including ours, indicate that this gap can be narrowed for specific networks and tasks.
@@ Appendix B Model hyperparameters and training details
All hyperparameters were carefully tuned per method and problem to ensure fair comparisons across methods. The number of episodes was chosen to illustrate the convergence of our method relative to baseline methods, with 8000, 20000, and 1000 episodes used for the Acrobot, Cartpole, and Reaching problems, respectively. In all our experiments, we set the smoothing factor for reward estimation to $\lambda=0.66$, striking a balance between recent values (for quick adaptability) and a longer history (for robustness against rapid reward fluctuations). Each neural network consisted of an input, hidden, and output layer. Input units corresponded to environment observation elements, and output units to possible actions. Specifically, Acrobot, Cartpole, and Reaching used 6, 4, and 32 input units and 3, 2, and 2 output units, respectively, with hidden layer sizes of 64, 64, and 128. Table 2 summarizes the learning rate $\eta$ and noise standard deviation $\sigma$ for each method and problem. Higher values of these parameters could lead to unstable training, while lower values may result in slower learning.

### v4

@@ 1 Introduction
Backpropagation (BP) remains the workhorse algorithm for training artificial neural networks, computing precise global gradients for weight updates [54]. However, its operational requirements are fundamentally incompatible with the constraints of neuromorphic hardware [10]. BP requires precise global error propagation, which translates into high-bandwidth global data movement and large memory requirements to store activations, the main source of energy consumption in deep learning systems [41]. This contrasts sharply with the hardware needs for local synaptic updates that rely only on information available at the synaptic weight. In addition, BP’s iterative forward and backward passes are incompatible with the asynchronous event-driven regime in which most neuromorphic devices operate [47], and its deterministic nature cannot tolerate inherent noise in neuromorphic circuits.
In response, local plasticity approaches and biologically-inspired learning rules have been proposed to bypass the constraints of BP [73, 72, 35, 55, 25, 76]. A particularly promising family of methods is noise-based learning, which leverages stochastic perturbations (either in weights [12, 78, 5, 9] or activations [19, 24, 74, 71, 15]) to approximate the true direction of the gradient. By injecting small random perturbations, these methods avoid the need for perfectly matched feedback pathways and can leverage inherent noise [60, 40].
Given the prevalence of noise on neuromorphic and biological substrates [11, 62, 42], using noise as a mechanism to learn synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [36, 43] as a promising candidate. RMHL, which is a form of three-factor Hebbian learning [21, 33], offers a potentially powerful mechanism for credit assignment without explicit backpropagation of errors. Despite theoretical advances, few noise-based methods have been adapted to real-world tasks. Some attempts which have been made include application to control problems [4], for spatiotemporal pattern generation [29], and for training non-differentiable spiking neural networks [14], and these particularly struggle with delayed reward regimes. This gap has limited the integration of noise-based approaches for effective credit assignment in temporally extended setups.
@@ 2.2.2 Noise-based learning rule
where $K$ denotes the number of samples and $\tilde{\theta}_{t}^{(i)}=\theta_{t}+\epsilon^{(i)}_{t}$ the noise-perturbed parameters (see Appendix A).
In practice, we consider $K=1$, analogous to single-sample updates in stochastic gradient descent. Although increasing $K$ (using multiple noise samples per step) would reduce the variance of the gradient estimate, it would linearly increase the computational cost and latency by requiring $K$ forward passes per update. Here, $g\bigl(\tilde{\theta}^{(i)}_{t}\bigr)=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})$ and $g\left(\theta_{t}\right)=\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$ are the log of the noise-perturbed and noise-free output, respectively. Thus, we define $\rho_{t}=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})-\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$, which captures the impact of the noise on the policy. Computationally, this term measures the influence of the perturbation on the action selection, as it quantifies how much the injected noise increased or decreased the likelihood of the chosen action compared to the noise-free baseline. Putting this together for the $K=1$ case, we obtain our noise-based learning rule
@@ 2.2.3 Learning with node level noise in neural network
To formulate the learning rule for specific layer parameters $W^{l}_{t}\in\{W^{1}_{t},\ldots,W^{L}_{t}\}$, we re-express the gradient in terms of the pre-activations instead of the parameters themselves. Here, the general parameter vector $\theta_{t}=\{W^{1}_{t},\ldots,W^{L}_{t}\}$ corresponds to the full set of all weight matrices for all layers at time $t$. To do so, we first rewrite the gradient term as $\nabla g\left(W^{l}_{t}\right)=\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)$ and apply the chain rule to break down the gradient with respect to $W^{l}_{t}$ as
@@
where $i$ indexes over a set of repeated samples of the noise term $\xi$. Again, as above, by reducing this to the single sample ($n=1$) case, we can define a layer-specific weight update rule
@@ 2.2.4 Learning from sparse rewards
Our derivations so far assume that synaptic updates happen at every time step, implying a continuous stream of rewards and learning signals. In reality, however, rewards are often sparse, arriving only after a sequence of actions or upon reaching specific milestones. To handle this, we will modify our learning rule to account for rewards received at arbitrary times.
Let us denote these discrete reward times as $\tau_{m}\in\{\tau_{0},\tau_{1},\dots,\tau_{M}\}$. At these moments, we compute the RPE as
@@ 2.4 Experimental validation
We compare NRL to two baselines: an exact-gradient version of NRL, which serves as an “optimal performance” benchmark, and an RMHL approach.
@@
Table 1: Learning rules. The terms in the BPTT update are rearranged for easier comparison. Updates on NRL and RMHL rely on local synaptic information (inputs $\tilde{x}$, noise $\xi$) modulated by global broadcasted scalars (RPE $\delta$, and for NRL, noise impact $\rho$). In contrast, BPTT relies on non-local gradient information ($\nabla$) propagated from later layers. The updates are given by $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\Delta W^{l}_{\tau_{m}}$.
@@
NRL $\Delta W^{l}_{\tau_{m}}=\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{\,l}_{t}\,\rho_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$
BPTT $\Delta W^{l}_{\tau_{m}}=\delta_{\tau_{m}}\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})$
RMHL $\Delta W^{l}_{\tau_{m}}=\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{\,l}_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$
In environments with high reward variability, like Cartpole and Acrobot, we stabilize synaptic updates by scaling the RPE by dividing by $r_{\tau_{m}}$. This normalization accounts for the non-stationary nature of the return in some tasks, where the reward magnitude scales with trial duration. By treating the error as relative to the current return, we prevent update instability as the agent’s performance improves. Hyperparameter values and training details are provided in Appendix B. All experiments are implemented in Python using PyTorch [50]. Our models and scripts are available for reproducibility at https://github.com/jesusgf96/noise-based-reward-modulated-learning.
@@ 3.1 Performance on control tasks with instantaneous and delayed rewards
To ensure comparability with RMHL methods, which typically utilize single-hidden-layer networks, we first conduct experiments with a one-hidden-layer neural network for BP, NRL, and RMHL. Training details are provided in Appendix B. In Section “Scalability to deeper architectures”, we extend the comparison to deeper networks with multiple hidden layers.
@@ 3.3 Learning using only noisy passes
To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem. We chose this task as it offers a representative delayed-reward challenge, unlike the Reaching task. Its stable dynamics also allowed us to clearly separate the approximation error from the variance inherent to the task itself, which was difficult with Cartpole’s sensitive initial conditions. The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Fig. 4A. Furthermore, Fig. 4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with a minimum of 2 noisy passes. We also extended this evaluation to 10 noisy passes to explore the impact of increasing the number of passes, showing slightly faster initial convergence.
@@ 4 Discussion
Our results demonstrate the learning effectiveness of NRL, solving both immediate and delayed feedback/reward problems in simulated environments, achieving performance comparable to baselines optimized employing backpropagation. In addition, it shows a significant improvement over RMHL [36, 43], a modern approach that also uses noise as the core of its synaptic plasticity. Our scalability results further demonstrate that NRL effectively assigns credit even in multilayer networks, a challenge for many biologically plausible or local algorithms, indicating its potential for scaling. This is a critical difference from RMHL, which struggles to assign credit in multilayer networks, as evidenced by its minimal performance improvements across trials. Nevertheless, as the depth of the network increases, NRL shows slower convergence due to its noise-driven updates. Beyond network depth, NRL performance is also influenced by the duration of the reward delay. As this delay increases, the eligibility trace accumulates more noise, lowering the signal-to-noise ratio in the gradient estimate. Consequently, although our RPE normalization prevents instability caused by large reward values in long episodes, we expect convergence to slow down as the delay extends, reflecting the difficulty of assigning credit over long time windows. This challenge can be compared with the vanishing gradient problem in BPTT, characterized by a signal decay. Yet, it contrasts with the additive variance accumulation problem seen in NRL, where the learning signal is not lost, but rather obscured by the noise of non-causal perturbations.
Regarding the scope of applicability, our framework suggests specific behaviors in complex scenarios. For high-dimensional observation spaces, the main challenge for perturbation-based learning is the potential variance increase in the gradient estimate. In this scenario, NRL may require more samples for convergence. However, the inclusion of a baseline via the RPE serves as a variance reduction mechanism [74], stabilizing the learning signal even in larger input spaces. When it comes to handling continuous action spaces, NRL, as a policy gradient approximation, is naturally compatible with continuous control. This extension would not require any changes in the learning rule, but simply replacing the categorical output with a Gaussian policy head, where noise perturbs the action mean directly.
NRL shares similarities with node perturbation methods [19, 24, 74, 71, 15], which also use noise to guide learning. However, they differ in their optimization objective and temporal scope. Standard node perturbation is formulated as a zeroth-order optimization technique for estimating the gradient of an immediate loss function. In contrast, NRL is derived as an approximation of policy gradients, designed to maximize long-term expected return. This distinction is realized through the use of an eligibility trace [23, 26], which keeps a temporary memory of local information (perturbation, activations, etc) until a delayed reward arrives. This effectively assigns credit to past neural states when actions and outcomes are separated in time. In addition, NRL also incorporates variance reduction via the RPE ($\delta_{t}=r_{t}-\bar{r}_{t}$), where the expected reward $\bar{r}_{t}$ acts as a baseline to stabilize the gradient estimate.
@@
Additionally, the local-update structure of NRL provides direct compatibility with neuromorphic substrates [59], where only locally available information in the synapse is prioritized and global communication is severely penalized. This eliminates the need for the high-bandwidth global data movement and large memory buffers required by backpropagation, the primary source of energy consumption in deep learning models, addressing the fundamental energy and latency bottlenecks of conventional Von Neumann systems. Furthermore, NRL’s reliance on noise for gradient approximation offers a paradigm shift: it transforms the device stochasticity common in neuromorphic circuits from an engineering challenge into a computational resource. The inherent noise in NRL is also advantageous for reinforcement learning setups, as it naturally introduces the stochasticity required for exploration [28, 34]. Injected noise perturbs neural states, promoting deviations from expected values and thus introducing variations in the agent’s policy. Recent studies on noise-based adaptation [12] have demonstrated that noise-driven learning mechanisms can effectively balance exploration and exploitation, particularly in volatile environments.
Despite its promise, our work has limitations that suggest future research directions. The current requirement for both noisy and clean forward passes (or averaging multiple noisy passes) per synaptic update, while theoretically simple, introduces a time-latency and computational overhead. In practice, this overhead is mitigated by the accelerated dynamics of neuromorphic hardware, which typically operates at microsecond timescales, orders of magnitude faster than real-world environments [8]. Thus, performing multiple forward passes potentially has a negligible impact on latency. From an energy perspective, the cost of extra passes is offset by eliminating the high-bandwidth data transport and memory buffering required by gradient-based methods, such as BP [13], which remain the primary energy bottlenecks in conventional learning systems. However, to further enhance this method, future work could explore alternative ways to estimate noise impacts, eliminating the need for multiple passes.
In addition, while we focused on a rate-based model, the deployment of NRL on physical hardware like Intel Loihi [8] requires its extension to spiking neural networks (SNNs). Adapting NRL to this domain presents specific implementation challenges. For instance, the continuous eligibility traces used here would need to be replaced by synaptic tracers that accumulate discrete spike events via exponential decay kernels [26]. Similarly, the injected noise used for gradient approximation could translate into stochastic membrane potential fluctuations or probabilistic firing thresholds [40], transforming perturbations into spike-timing variability. Also, while our derivation assumes Gaussian noise, physical hardware often exhibits non-Gaussian stochasticity. However, as elaborated in Appendix C, the central limit theorem suggests that the aggregation of these discrete noise sources at the neuronal level effectively approaches a Gaussian distribution. Thus, our theoretical framework is expected to remain robust in physical neuromorphic systems.
@@
Furthermore, the simplicity of our RPE calculation limits its robustness in more complex tasks requiring long-term planning. Future development should incorporate more sophisticated temporal difference learning mechanisms [63, 65], such as a learned value function within an actor-critic architecture [31], to provide a richer temporally sensitive estimate of rewards. Finally, while backpropagation may offer faster convergence in unconstrained settings, recent studies [7, 13], including ours, indicate that this gap can be narrowed for specific networks and tasks.
Finally, regarding sample efficiency, NRL is an on-policy algorithm theoretically grounded in the policy gradient theorem [66]. Although off-policy methods can improve data efficiency, they typically rely on mechanisms that conflict with biological and neuromorphic constraints. Incorporating off-policy mechanisms, such as experience replay [44] or importance sampling [52], would require centralized memory buffers and global data transport, violating the principle of synaptic locality. Crucially, this restriction allows us to isolate and evaluate the performance of the learning rule itself, avoiding the obscuring factors introduced by add-ons. Future work could investigate how such off-policy augmentations can be adapted to the neuromorphic context to further enhance sample efficiency.
@@ Appendix B Model hyperparameters and training details
All hyperparameters were carefully tuned per method and problem to ensure fair comparisons across methods. The number of episodes was chosen to illustrate the convergence of our method relative to baseline methods, with 8000, 20000, and 1000 episodes used for the Acrobot, Cartpole, and Reaching problems, respectively.
In all our experiments, we set the smoothing factor for reward estimation to $\lambda=0.66$. This value was empirically found to offer a robust trade-off across all tasks, striking a balance between recent values (for quick adaptability) and a longer history (for robustness against rapid reward fluctuations). We observed that this choice was sufficiently efficient to be applied uniformly across all environments without the need for task-specific adjustment.
Each neural network consisted of an input, hidden, and output layer. Input units corresponded to environment observation elements, and output units to possible actions. Specifically, Acrobot, Cartpole, and Reaching used 6, 4, and 32 input units and 3, 2, and 2 output units, respectively, with hidden layer sizes of 64, 64, and 128. Table 2 summarizes the learning rate $\eta$ and noise standard deviation $\sigma$ for each method and problem. Higher values of these parameters could lead to unstable training, while lower values may result in slower learning.
@@
Appendix C Noise distributions in neuromorphic hardware
In our derivation of the NRL update rule, we assume that the perturbative noise $\xi$ follows a Gaussian distribution. This assumption is crucial for maintaining equality between the expected update and the true gradient. However, in physical neuromorphic implementations, noise sources at the device level may follow different distributions.
In practice, neuromorphic substrates manifest stochasticity through physical processes. Thermal or Johnson–Nyquist noise [27], arising from charge carriers, is inherently Gaussian and is always present in analog circuits. However, other sources are non-Gaussian. Shot noise [56], caused by the discrete nature of electric charge crossing a barrier, follows a Poisson distribution. Random telegraph noise [30], common in memristive devices, manifests as discrete switching between two states, following a bimodal distribution.
Although individual synaptic or neuronal noise sources may follow non-Gaussian distributions, the NRL update rule operates on the aggregated activity influencing the neuron’s state. A fundamental property of neural integration is the summation of inputs from multiple presynaptic sources. According to the central limit theorem (CLT) [17], the sum of a large number of independent random variables, regardless of their original distribution, provided they have finite variance, converges to a Gaussian distribution. Consequently, the Gaussian assumption in $\xi$ used in our derivation is not only a mathematical convenience, but a statistically valid approximation of the effective aggregate noise seen in hardware implementations or biological networks.
Even in scenarios with low connectivity where the CLT approximation is not perfect, the learning rule is expected to remain robust. While Theorem 25 relies on this Gaussian assumption for exact equality between the expected update and the true gradient, in practice, optimization does not require the update vector to be identical to the true gradient, but to have a positive projection onto it, i.e., a positive cosine similarity. This principle is well-established in the literature, where algorithms, such as feedback alignment [38], direct feedback alignment [48], or sign-symmetry [37], demonstrate that approximate gradients are sufficient for convergence, as long as they remain within $90^{\circ}$ of the steepest descent direction. Therefore, skewed non-Gaussian noise may reduce convergence speed by lowering this cosine similarity, but it does not orthogonalize the update vector, thus maintaining a valid descent trajectory.